Your grandmother invested $2,300 for you on the day you were born. This investment has earned an average of 13.8 percent annually. How old are you if the investment is now worth $18,642? O 14.36 years 14.58 years 15.15 years 16.19 years
You just inherited $7100 which you are investing at 7.25 percent interest. How long will you have to wait un your account reaches a value of $12,000 O 6.96 years O 8.34 years O 7.50 years O 8.55 years
First City Bank pays 7 percent simple interest on its savings account balances, whereas Second City Bank pays 7 percent interest compounded annually. If you made a $73,000 deposit in each bank, how much more money would you earn from your Second City Bank account at the end of 9 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16 .) Difference in accounts
Six years ago, you convinced your son to deposit $900 in a savings account at the bank that will pay 5.25 percent interest. Currently, he is itching to spend that money. How much more money will he have if you can convince him to wait another year to withdraw his savings rather than withdrawing the funds today? O $55.09 $69.70 $49.73 O $64.23
Answer #1

1.

Grandmother Invested = $ 2,300

Current Value of Investment = $ 18,642

Interest Rate = 13.8 % annually

If Interest compounded annually, then -

Future Value of Initial Investment = Initial Investment(1 + i )t

Where,

i = Interest rate

t = Time (years)

In above case, Initial Investment = $ 2,300

Future Value of Initial Investment in t(time) = $ 18,642

Now, Computation of Time(t)

2,300(1 + 0.138)t = 18,642

(1.138)t = 18,642/2,300

(1.138)t = 8.1052

Applying Log both side,

t*Log(1.138) = Log(8.1052)

t = Log(8.1052) / Log (1.138)

t = 0.9088 / 0.0561

t = 16.19 years

It takes 16.19 years at 13.8 % annually to make initial investment $ 2,300, worth $ 18,642

2.

Initial Investment = $ 7,100

Future value of Investment = $ 12,000

Interest rate = 7.25 % or 0.0725

Time = t

Computation of Time(t)

7,100(1 + 0.0725)t = 12,000

(1.0725)t = 12,000/7,100

(1.0725)t = 1.690

Applying Log both side,

t*Log(1.0725) = Log(1.690)

t = Log(1.690) / Log (1.0725)

t = 0.2279 / 0.0304

t = 7.50 years

It takes 7.50 years at 7.25 % annually to make initial investment $ 7,100, worth $ 12,000

3.

Simple Interest method refers where interest is fixed for each period and Compound Interest method refers where Interest calculated on compound value (Principal + Previous Interest) for each period.

Simple Interest Method

A = P + P*R*T

Where,

A = Amount (Future Value)

P = Principal (Initial Investment)

R = Interest Rate

T = Time

Compound Interest Method

A = P(1+R)T

Where,

A = Amount (Future Value)

P = Principal (Initial Investment)

R = Interest Rate

T = Time

In above case, $ 73,000 in deposited in each Bank

First city Bank pays 7% simple interest for 9 years

Therefore, Deposit value after 9 years would be

A = P + P*R*T

A = 73,000 + 73,000*0.07*9

A = 73,000 + 45,990

A = 118,990

Deposit value after 9 years would be $ 118,990 in First City Bank.

Second city Bank pays 7% Compound interest for 9 years

Therefore, Deposit value after 9 years would be

A = P(1+R)T

A = 73,000(1.07)9

A = 73,000(1.8385)

A = 134,208.52

Deposit value after 9 years would be $ 134,208.52 in Second City Bank.

Therefore, Difference in Account = $ 134,208.52 - $ 118,990

= $ 15,218.51

4.

Deposit (P) = $ 900

Interest Rate (R) = 0.0525

Calculation of Amount (A) after 6 years if interest compounded annually.

A = P(1+R)T

A = 900(1.0525)6

A = 900(1.35935)

A = 1,223.42

Calculation of Amount (A) after 6 years if interest compounded annually.

A = P(1+R)T

A = 900(1.0525)7

A = 900(1.43072)

A = 1,287.65

He will withdraw extra money amounting $ 64.23 ( $ 1.287.65 - $ 1,223.42) if he wait for one more years i.e end of 7th year.